# Jet definition and TMD factorisation in SIDIS

### v1

@@ Introduction.
The first term in the square bracket in Eq. (12) is the Sudakov double logarithm that we heuristically derived in Eq. (4). It agrees with the expectation from the CSS kernel for the quark TMD alone Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). The second term is a Sudakov single logarithm, which depends both on the quark anomalous dimension $\Gamma_{q}=\frac{3\alpha_{s}C_{F}}{4\pi}$ and on the jet parameter $R$. By only keeping these logarithmic terms, one obtains the virtual contribution to the DGLAP+CSS evolution of the quark TMD, conveniently rewritten as
$\left.x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\right|_{V}=-\frac{\alpha_{s}C_{F}}{2\pi}\,x\mathcal{F}_{q}(x,\bm{P}_{\perp})$
×$\quad\qquad\qquad\qquad\times\int_{P_{\perp}^{2}}^{Q^{2}}\frac{\mathrm{d}\ell_{\perp}^{2}}{\ell_{\perp}^{2}}\left[\ln\left(\frac{Q^{2}}{R^{2}P_{\perp}^{2}}\right)-\frac{3}{2}\right],$ (13)
@@
×$\times\!\int_{x_{\star}}^{1}{\mathrm{d}\xi}\,\mathcal{P}_{qq}(\xi)\,\frac{x}{\xi}f_{q}\left(\frac{x}{\xi},P_{\perp}^{2}\right),$ (14)
where we have also inserted a running coupling, as standard in this context. The CSS equation obeyed by the quark TMD can be simply obtained by taking a derivative w.r.t. $\ln Q^{2}$ in the sum of Eqs. (11) and (13).
To summarise, we have demonstrated that TMD factorisation for jet production in SIDIS is not guaranteed by all jet definitions and we have designed a longitudinally-invariant jet clustering algorithm which preserves both the factorisation and the universality of the DGLAP+CSS evolution of the quark TMD at small $x$. Physically, this jet definition is able to resolve the DGLAP and Sudakov dynamics of the struck sea quark. While these results are derived in the context of the high energy factorisation, they also apply at moderate values of $x$, as demonstrated in the Supplemental Material, where we find a similar factorisation property by following the dynamics of the struck quark in the target picture. Once again, the proper choice of the jet definition turns out to be essential to ensure the correct matching between the NLO corrections to SIDIS and the DGLAP+CSS evolution of the quark TMD. It will be important to investigate further the jet definition Eq. (5) both at small and moderate $x$, as we anticipate that it could be of great importance for the forthcoming quark tomography at the EIC with jets Arratia et al. (2020a); Page et al. (2020).
@@ Appendix A Supplemental material
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}\times x\mathcal{F}_{q}(x,\bm{P}_{\perp})\left[1+\mathcal{O}\left(\frac{P_{\perp}^{2}}{Q^{2}}\right)\right]$ (42)
@@
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})\equiv\frac{N_{c}}{\pi^{2}}\int\mathrm{d}^{2}\bm{b}_{\perp}\int\frac{\mathrm{d}^{2}\bm{q}_{\perp}}{(2\pi)^{2}}\ \mathcal{D}(x,\bm{q}_{\perp})\left[1-\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{(P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2})}\,\ln\frac{P_{\perp}^{2}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\right]$ (43)
@@
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})$ $=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\frac{1}{3}\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}(x,\bm{q}_{\perp})\,,\quad x\mathcal{G}_{D}(x,\bm{q}_{\perp})=\frac{N_{c}\bm{q}_{\perp}^{2}}{2\pi^{2}\alpha_{s}}\int\mathrm{d}^{2}\bm{b}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})$ (44)
where we have recognised the so-called gluon dipole TMD $x\mathcal{G}_{D}(x,\bm{q}_{\perp})$ Dominguez et al. (2011). The $1/3$ coefficient physically comes from the integral of the $g\to q\bar{q}$ splitting function. Indeed, Eq. (44) can be equivalently rewritten, in the small $x$ regime, in term of the gluon PDF $xG(x,P_{\perp}^{2})$:
$x\mathcal{F}_{q}(x,\bm{P}_{\perp})=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{0}^{1}\mathrm{d}\xi\ P_{qg}(\xi)\frac{x}{\xi}G\left(\frac{x}{\xi},P_{\perp}^{2}\right)\,,\quad xG(x,P_{\perp}^{2})=\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}(x,\bm{q}_{\perp}).$ (45)
@@
Let us now turn to the differences between the two classes of jet definitions, $A$ and $B$, as manifest from the inspection of the Sudakov logarithms in Eqs. (107) and (108). Clearly, the main difference refers to the coefficient of the Sudakov double logarithm. With the clustering condition $B$, this coefficient is equal to $-\alpha_{s}C_{F}/(4\pi)$ (like in Eq. (12) in the Letter) and is consistent with TMD factorisation, as also explained in the Letter: (i) it has the right value to cancel the (single) real Sudakov log in Eq. (11) of the Letter after integrating over $P_{\perp}$, and (ii) it is furthermore consistent with expectations from the CSS formalism Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). On the contrary, the respective coefficient predicted by the clustering condition $B$, that is, $-3\alpha_{s}C_{F}/(4\pi)$, not only is inconsistent with the general CSS results aforementioned, but also does not fulfil the condition of real–virtual cancellation after integrating over $P_{\perp}$; indeed, with this jet definition, the real Sudakov logarithm is found to be twice as large as that shown in the second line of Eq. (11), hence it cannot cancel against a virtual double logarithm which is three times larger than that in Eq. (108).
@@
$\mathcal{F}_{q}^{\rm\,(unsub.)}(x,\bm{P}_{\perp})$ $=$ $\frac{1}{2}\int\frac{\mathrm{d}\xi^{-}\mathrm{d}^{2}\bm{\xi}_{\perp}}{(2\pi)^{3}}e^{-ix\xi^{-}P^{+}+i\bm{\xi}_{\perp}\cdot\bm{P}_{\perp}}\left\langle P\left|\overline{\psi}(\xi){\cal L}_{n}^{\dagger}(\xi)\gamma^{+}{\cal L}_{n}(0)\psi(0)\right|P\right\rangle\ ,$ (126)
@@
Here, $\bm{b}_{\perp}$ is the Fourier conjugate variable with respect to the transverse momentum $\bm{P}_{\perp}$ and $\mu_{F}$ is the factorization scale. And $\zeta_{c}^{2}$ is defined as $\zeta_{c}^{2}=x^{2}(2v\cdot P)^{2}/v^{2}=2(xP^{+})^{2}e^{-2y_{n}}$ with $y_{n}$ being the rapidity cut-off in the Collins-11 scheme. The second factor represents the soft factor subtraction with $n$ and $\bar{n}$ as the light-front vectors $n=(1^{-},0^{+},0_{\perp})$, $\bar{n}=(0^{-},1^{+},0_{\perp})$, whereas $v$ is an off-light-front $v=(v^{-},v^{+},0_{\perp})$ with $v^{-}\gg v^{+}$.
@@
where $\mu_{F}$ is the factorisation scale and $\zeta_{c}^{2}\equiv x^{2}(2v\cdot P_{N})^{2}/v^{2}=2(xP^{-}_{N})^{2}e^{-2y_{n}}$ with $y_{n}$ being the rapidity cut-off in the Collins-11 scheme ($v^{\mu}=(v^{-},v^{+},0_{\perp})$ is a slightly off-light-front 4-vector with $v^{+}\gg v^{-}$).
@@
To conclude this discussion, let us recall that the NLO corrections to jet production in the valence quark picture are unable to distinguish between different values of the parameter $\beta$ in the $B$ class of clustering conditions, Eq. (32). The $\beta$–dependence should however become visible at NNLO, or two loop, order. Our previous calculation using sea quarks in the dipole picture predicts that the only acceptable value (in the sense of the conformity with TMD factorisation) is $\beta=0$. If the DIS jets are clustered in the Breit frame by using spherically invariant distance measures as in Refs. Gutierrez-Reyes et al. (2018, 2019a, 2019b) (which correspond to Eq. (32) with $\beta=2$), we therefore expect violations of TMD factorisation starting with the two loops order.

### v2

@@ Introduction.
where $x_{\star}$ is a $\mathcal{O}(1)$ number that marks the separation between the phase spaces $z_{g}\leq x_{\star}k_{g\perp}^{2}/Q^{2}$ contributing to collinearly improved BK/JIMWLK evolution Beuf (2014); Iancu et al. (2015); Hatta and Iancu (2016); Ducloué et al. (2019) of the quark TMD, and $z_{g}\geq x_{\star}k_{g\perp}^{2}/Q^{2}$ contributing to the NLO impact factor. Since the latter has undergone a power expansion in $P_{\perp}/Q$ and $Q_{s}/P_{\perp}$, unlike the former, the cancellation of the $x_{\star}$ dependence between small $x$ evolution and NLO impact factor is not complete anymore, but the remaining $x_{\star}$ dependence is a pure NLO effect, as clear from Eq. (12).
The first term in the square bracket in Eq. (12) is the Sudakov double logarithm that we heuristically derived in Eq. (4). It agrees with the expectation from the CSS kernel for the quark TMD alone Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). The second term is a Sudakov single logarithm, which depends both on the quark anomalous dimension $\Gamma_{q}=\frac{3\alpha_{s}C_{F}}{4\pi}$ and on the jet parameter $R$. By only keeping these logarithmic terms, one obtains the virtual contribution to the DGLAP+CSS evolution of the quark TMD:
$\left.x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\right|_{V}=-\frac{\alpha_{s}C_{F}}{2\pi}\,x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})$
×$\quad\quad\times\left[\frac{1}{2}\ln^{2}\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)-\left(\frac{3}{2}+\ln(R^{2})\right)\ln\left(\frac{Q^{2}}{P_{\perp}^{2}}\right)\right].$ (13)
@@
×$\times\!\int_{x}^{1}{\mathrm{d}\xi}\,\mathcal{P}_{qq}(\xi)\,\frac{x}{\xi}f_{q}\left(\frac{x}{\xi},P_{\perp}^{2}\right),$ (14)
where we have also inserted a running coupling, as standard in this context. The CSS equation obeyed by the quark TMD can be obtained by taking a derivative w.r.t. $\ln Q^{2}$ in the sum of Eqs. (11) and (13). The precise relation between our “top-down” approach Caucal and Iancu (2024) to the resummation of Sudakov logarithms and the standard CSS formalism is given by $x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})=x\mathcal{F}_{q}^{\rm(sub)}(x,\bm{P}_{\perp},\mu_{F}=Q,\zeta_{c}=Q)$ where $\mathcal{F}_{q}^{\rm(sub)}$ is the subtracted quark TMD in Collins-11 scheme Collins (2023) and $\mu_{F}$, $\zeta_{c}$ are respectively the UV and rapidity renormalisation scales, both identified with the hard scale $Q$ of the process  [^6] For $\zeta_{c}=Q$, our longitudinal upper cut-off in $\xi\leq 1-P_{\perp}/Q$ introduced the jet constraint is equivalent to the rapidity cut-off $y_{n}$ in Collins-11 scheme defined as $\zeta_{c}^{2}=2(x_{\rm Bj}P_{N}^{-})e^{2y_{n}}$.. As shown in Ebert et al. (2022); del Rio et al. (2024), this diagonal scheme also has the advantage of preserving the validity of Eq. (9) connecting the quark TMD to the quark PDF (with $x\mathcal{F}_{q}^{(0)}\to x\mathcal{F}_{q}(x,\bm{\ell}_{\perp},P_{\perp}^{2})$) up to $\mathcal{O}(\alpha_{s}^{2})$ corrections.
In the end, the single inclusive jet cross-section at small $x$ and $Q^{2}\gg P_{\perp}^{2}\gg Q_{s}^{2}$, for our new jet distance measure, can be written
$\left.\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm NLO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})$
×$\times\left[1-\frac{3\alpha_{s}C_{F}}{2\pi}\ln(R)+\mathcal{O}(\alpha_{s})\right]\,,$ (15)
where all the potentially large logarithms $\alpha_{s}\ln(1/x)$, $\alpha_{s}\ln(P_{\perp}^{2}/\Lambda^{2})$, $\alpha_{s}\ln^{2}(Q^{2}/P_{\perp}^{2})$ and $\alpha_{s}\ln(Q^{2}/P_{\perp}^{2})$ are resummed within the quark TMD via BK/JIMWLK — or Balitsky-Fadin-Kuraev-Lipatov (BFKL) Kuraev et al. (1977); Balitsky and Lipatov (1978) in the dilute limit $P_{\perp}\gg Q_{s}$ — and DGLAP+CSS evolution equations, respectively.
To summarise, we have demonstrated that TMD factorisation for jet production in SIDIS is not guaranteed by all jet definitions and we have designed a longitudinally-invariant jet clustering algorithm which preserves both the factorisation and the universality of the DGLAP+CSS evolution of the quark TMD at small $x$. Physically, this jet definition is able to resolve the DGLAP and Sudakov dynamics of the struck sea quark and to distinguish the backward antiquark jet (in the Breit frame) from the beam remnant. While these results are derived in the context of the high energy factorisation, they also apply at moderate values of $x$, as demonstrated in the Supplemental Material, where we find a similar factorisation property by following the dynamics of the struck quark in the target picture. Once again, the proper choice of the jet definition turns out to be essential to ensure the correct matching between the NLO corrections to SIDIS and the DGLAP+CSS evolution of the quark TMD. It will be important to investigate further the jet definition Eq. (5) both at small and moderate $x$, as we anticipate that it could be of great importance for the forthcoming quark tomography at the EIC with jets Arratia et al. (2020a); Page et al. (2020).
@@ Appendix A Supplemental material
$\left.\frac{\mathrm{d}\sigma^{\gamma_{\rm T}^{\star}+A\to q+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\right|_{\rm LO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}\times x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})\left[1+\mathcal{O}\left(\frac{P_{\perp}^{2}}{Q^{2}}\right)\right]$ (43)
@@
$x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})\equiv\frac{N_{c}}{\pi^{2}}\int\mathrm{d}^{2}\bm{b}_{\perp}\int\frac{\mathrm{d}^{2}\bm{q}_{\perp}}{(2\pi)^{2}}\ \mathcal{D}(x,\bm{q}_{\perp})\left[1-\frac{\bm{P}_{\perp}\cdot(\bm{P}_{\perp}-\bm{q}_{\perp})}{(P_{\perp}^{2}-(\bm{P}_{\perp}-\bm{q}_{\perp})^{2})}\,\ln\frac{P_{\perp}^{2}}{(\bm{P}_{\perp}-\bm{q}_{\perp})^{2}}\right]$ (44)
@@
$x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})$ $=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\frac{1}{3}\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}^{(0)}(x,\bm{q}_{\perp})\,,\quad x\mathcal{G}_{D}^{(0)}(x,\bm{q}_{\perp})=\frac{N_{c}\bm{q}_{\perp}^{2}}{2\pi^{2}\alpha_{s}}\int\mathrm{d}^{2}\bm{b}_{\perp}\mathcal{D}(x,\bm{q}_{\perp})$ (45)
where we have recognised the so-called gluon dipole TMD $x\mathcal{G}_{D}^{(0)}(x,\bm{q}_{\perp})$ Dominguez et al. (2011). The $1/3$ coefficient physically comes from the integral of the $g\to q\bar{q}$ splitting function. Indeed, Eq. (45) can be equivalently rewritten, in the small $x$ regime, in term of the gluon PDF $xG(x,P_{\perp}^{2})$:
$x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})=\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\int_{0}^{1}\mathrm{d}\xi\ P_{qg}(\xi)\frac{x}{\xi}G\left(\frac{x}{\xi},P_{\perp}^{2}\right)\,,\quad xG^{(0)}(x,P_{\perp}^{2})=\int_{0}^{P_{\perp}^{2}}\mathrm{d}^{2}\bm{q}_{\perp}\ x\mathcal{G}_{D}^{(0)}(x,\bm{q}_{\perp}).$ (46)
@@
Let us now turn to the differences between the two classes of jet definitions, $A$ and $B$, as manifest from the inspection of the Sudakov logarithms in Eqs. (108) and (109). Clearly, the main difference refers to the coefficient of the Sudakov double logarithm. With the clustering condition $B$, this coefficient is equal to $-\alpha_{s}C_{F}/(4\pi)$ (like in Eq. (12) in the Letter) and is consistent with TMD factorisation, as also explained in the Letter: (i) it has the right value to cancel the (single) real Sudakov log in Eq. (11) of the Letter after integrating over $P_{\perp}$, and (ii) it is furthermore consistent with expectations from the CSS formalism Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). On the contrary, the respective coefficient predicted by the clustering condition $A$, that is, $-3\alpha_{s}C_{F}/(4\pi)$, not only is inconsistent with the general CSS results aforementioned, but also does not fulfil the condition of real–virtual cancellation after integrating over $P_{\perp}$; indeed, with this jet definition, the real Sudakov logarithm is found to be twice as large as that shown in the second line of Eq. (11), hence it cannot cancel against a virtual double logarithm which is three times larger than that in Eq. (109).
@@
After adding together the virtual corrections as given by Eq. (109) with $\beta=0$ (equivalently, Eq. (13) in the Letter) and the real corrections shown in Eq. (11) from the Letter, we are finally in a position to exhibit the NLO result for the jet measurement in SIDIS in the approximations of interest. This is shown in Eq. (15) from the Letter, that we repeat here for convenience:
$\frac{\mathrm{d}\sigma^{\gamma^{*}_{\rm T}+A\to j+X}}{\mathrm{d}^{2}\bm{P}_{\perp}}\Big{|}_{\rm NLO}$ $=\frac{8\pi^{2}\alpha_{\rm em}e_{f}^{2}}{Q^{2}}x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})\left[1-\frac{3\alpha_{s}C_{F}}{2\pi}\ln(R)+\mathcal{O}(\alpha_{s})\right]\,,$ (110)
where it is understood that
$x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})=x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})+x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\Big{|}_{R}+x\mathcal{F}_{q}^{(1)}(x,\bm{P}_{\perp},Q^{2})\Big{|}_{V}\,.$ (111)
The (real and virtual) NLO corrections exhibit logarithmic dependences upon the hard scale $Q^{2}$, as visible in Eqs. (11) and (13). By taking a derivative w.r.t. $\ln Q^{2}$, one can promote these one-loop corrections to an evolution equation, which reads
$\frac{\partial\,x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})}{\partial\ln Q^{2}}=$ $\,\frac{C_{F}}{2\pi}\left\{\frac{\alpha_{s}(P_{\perp}^{2})}{P_{\perp}^{2}}\int^{P_{\perp}^{2}}_{\Lambda^{2}}{\rm d}\ell_{\perp}^{2}\,x\mathcal{F}_{q}(x,\bm{\ell}_{\perp},Q^{2})-\int_{P_{\perp}^{2}}^{Q^{2}}\frac{{\rm d}\ell_{\perp}^{2}}{\ell_{\perp}^{2}}\alpha_{s}(\ell_{\perp}^{2})\,x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})\right\}$
$+\frac{3}{2}\frac{\alpha_{s}(Q_{\perp}^{2})C_{F}}{\pi}\,x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})\,.$ (112)
In all the terms, $\alpha_{s}$ runs with the transverse momentum of the daughter gluons produced by the hard splitting.
As mentioned in the Letter, Eq. (C.5) can be recognised as the diagonal version of the Collins-Soper-Sterman (CSS) equations Collins and Soper (1981, 1982); Collins et al. (1985); Collins (2023). (The same equations have also been obtained with the SCET formalism Becher and Neubert (2011); Becher et al. (2013); Echevarria et al. (2012, 2014); Chiu et al. (2012); Ebert et al. (2020).) That is, this is the equation obeyed by the quark TMD $x\mathcal{F}_{q}^{\rm(sub)}(x,\bm{P}_{\perp},\mu_{F},\zeta_{c})$ provided the two resolution scales are identified with each other and with the hard scale of our problem, which is $Q$: $\mu_{F}=\zeta_{c}=Q$.
Eq. (C.5) should be solved as an initial value problem with the initial condition at $Q^{2}=P_{\perp}^{2}$ provided by Eq. (11) in the Letter, that is
$x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2}=P_{\perp}^{2})$ $\,=x\mathcal{F}_{q}^{(0)}(x,\bm{P}_{\perp})+\frac{\alpha_{s}}{2\pi^{2}}\frac{1}{P_{\perp}^{2}}\!\int_{x}^{1}\!{\mathrm{d}\xi}P_{qq}^{(+)}(\xi)\frac{x}{\xi}f_{q}\!\left(\frac{x}{\xi},P_{\perp}^{2}\right)$
$\,=\,\frac{1}{\pi}\frac{\partial xf_{q}(x,P_{\perp}^{2})}{\partial P_{\perp}^{2}}\,-\,\frac{3}{2}\frac{\alpha_{s}C_{F}}{2\pi^{2}}\,\frac{xf_{q}(x,P_{\perp}^{2})}{P_{\perp}^{2}}\,,$ (113)
where the second line follows after using the DGLAP equation, as shown in Eq. (14) in the Letter. Hence, in practice, one should first solve the DGLAP equation for the quark PDF $xf_{q}(x,P_{\perp}^{2})$ and then use its solution to build the initial condition for the CSS equation Eq. (C.5).
The impact-parameter version of the CSS equation is obtained after a Fourier transform from $\bm{P}_{\perp}$ to $\bm{b}_{\perp}$,
$x\tilde{\mathcal{F}}_{q}(x,\bm{b}_{\perp},Q^{2})\,\equiv\int\frac{{\rm d}^{2}\bm{P}_{\perp}}{(2\pi)^{2}}\,{\rm e}^{i\bm{P}_{\perp}\cdot\bm{b}_{\perp}}\,x\mathcal{F}_{q}(x,\bm{P}_{\perp},Q^{2})\,,$ (114)
and takes the expected from (the diagonal version of the CSS equations in $b_{\perp}$-space, see e.g. Eq. (4.12) in Boussarie et al. (2023))
$\frac{\partial x\tilde{\mathcal{F}}_{q}(x,\bm{b}_{\perp},Q^{2})}{\partial\ln Q^{2}}=$ $\frac{C_{F}}{\pi}\left[-\frac{1}{2}\int_{\mu_{b}^{2}}^{Q^{2}}\frac{{\rm d}\ell_{\perp}^{2}}{\ell_{\perp}^{2}}\,\alpha_{s}(\ell_{\perp}^{2})+\frac{3}{2}\alpha_{s}(Q^{2})\right]x\tilde{\mathcal{F}}_{q}(x,\bm{b}_{\perp},Q^{2}),$ (115)
where we have introduced the notation $\mu_{b}^{2}\equiv c^{2}_{0}/\bm{b}_{\perp}^{2}$ with $c_{0}=2\,e^{-\gamma_{E}}\simeq 1.123$ for the lower limit. Notice that the real term from the $P_{\perp}$-space equation Eq. (C.5) has (formally) disappeared after going to $b_{\perp}$-space: the Fourier transform has “corse-grained” the modes with $P_{\perp}\lesssim\mu_{b}$, for which real and virtual corrections have mutually cancelled. So, we are only left with the virtual corrections at $P_{\perp}>\mu_{b}$. This explains the lower limit $\mu_{b}^{2}$ in the above integral over $\ell_{\perp}^{2}$.
Eq. (115) can be easily solved for an arbitrary initial condition $x\tilde{\mathcal{F}}_{0}(x,\bm{b}_{\perp}^{2})\equiv x\tilde{\mathcal{F}}_{q}(x,\bm{b}_{\perp},Q^{2}=\mu_{b}^{2})$. However, the inverse Fourier transform to momentum-space is sensitive to very large values of $b_{\perp}$, for which a non-perturbative prescription is generally needed. Such prescriptions are well documented in the literature Collins (2023); Boussarie et al. (2023). Here, we would only like to add that the sensitivity to non-perturbative physics is alleviated by the phenomenon of gluon saturation: for sufficiently small values of $x$ and/or large values of the atomic number $A$, the saturation momentum is a semi-hard scale, $Q_{s}^{2}(x,A)\gg\Lambda^{2}$, and the initial condition $x\tilde{\mathcal{F}}_{0}(x,\bm{b}_{\perp}^{2})$ taken from the CGC effective theory, cf. Eq. (44), is rapidly decreasing at large distances, $b_{\perp}\gg 1/Q_{s}$, thus ensuring the convergence of the inverse Fourier transform.
@@
$\mathcal{F}_{q}^{\rm\,(unsub.)}(x,\bm{P}_{\perp})$ $=$ $\frac{1}{2}\int\frac{\mathrm{d}\xi^{+}\mathrm{d}^{2}\bm{\xi}_{\perp}}{(2\pi)^{3}}e^{-ix\xi^{+}P_{N}^{-}+i\bm{\xi}_{\perp}\cdot\bm{P}_{\perp}}\left\langle P\left|\overline{\psi}(\xi){\cal L}_{n}^{\dagger}(\xi)\gamma^{-}{\cal L}_{n}(0)\psi(0)\right|P\right\rangle\ ,$ (133)
@@
Here, $\bm{b}_{\perp}$ is the Fourier conjugate variable with respect to the transverse momentum $\bm{P}_{\perp}$ and $\mu_{F}$ is the factorization scale. And $\zeta_{c}^{2}$ is defined as $\zeta_{c}^{2}=x^{2}(2v\cdot P)^{2}/v^{2}=2(xP_{N}^{-})^{2}e^{2y_{n}}$ with $y_{n}$ being the rapidity cut-off in the Collins-11 scheme. The second factor represents the soft factor subtraction with $n$ and $\bar{n}$ as the light-front vectors $n=(0^{-},1^{+},0_{\perp})$, $\bar{n}=(1^{-},0^{+},0_{\perp})$, whereas $v$ is an off-light-front $v=(v^{-},v^{+},0_{\perp})$ with $v^{+}\gg v^{-}$.
@@
where $\mu_{F}$ is the factorisation scale and $\zeta_{c}^{2}$ has been defined above.
@@
To conclude this discussion, let us recall that the NLO corrections to jet production in the valence quark picture are unable to distinguish between different values of the parameter $\beta$ in the $B$ class of clustering conditions, Eq. (33). The $\beta$–dependence should however become visible at NNLO, or two loop, order. Our previous calculation using sea quarks in the dipole picture predicts that the only acceptable value (in the sense of the conformity with TMD factorisation) is $\beta=0$.